How would i solve these using factoring ? .

1- 5x^2 = 2x - 7x^2
2- 3x^2 + 6x = 2x^2 - 9
3- 5x^2 - 11x + 2 = 0

Answer by checkley71(8403)

5X^2=2X-7X^2
5X^2+7X^2-2X=0
12X^2-2X=0
2X(6X-1)=0
2X=0
X=0 ANSWER.
6X-1=0
6X=1
X=1/6 ANSWER.
---------------------------------------
3X^2+6X=2X^2-9
3X^2-2X^2+6X+9=0
X^2+6X+9=0
(X+3)(X+3)=0
X+3=0
X=-3 ANSWER.
------------------------------------
5X^2-11X+2=0
(5X-1)(X-2)=0
5X-1=0
5X=1
X=1/5 OR 0.2 ANSWER.
X-2=0
X=2 ANSWER.